OpenStudy (anonymous):

You save $15,000.00. You place one-third in a savings account earning a 4.6% APR compounded annually. You then invest one quarter of the remaining balance in a 3-year U.S. Treasury bond earning a 5.2% APR compounded annually and the rest in a stock plan. Your stock plan increases in value 3% the first year, decreases 8% in value the second year, and increases 6% in value the third year. What are the balances for each account by the end of the third year and the total gain on your original saved amount?

OpenStudy (anonymous):

S = Savings Account Final \[S = 15000/3 (1 + 0.046/1)^3\] = $5722.23 T = treasury bond final amount \[T = 10000/4(1+0.052/1)^3 = $2910.63\]

OpenStudy (anonymous):

Stock portion: $7500 to start and increases by 3% the first year. 7500 x 0.03 = 225 7500 + 225 = 7725 Second year it decreases by 8%: 7725 x .08 = 618 7725 - 618 = 7107 Increases by 6% the final year: 7107 x .06 = 426.42 7107+426.42 = 7533.42

OpenStudy (anonymous):

Savings Account = $5722.23 Treasury Bond = $2910.63 Stock = $7533.42

OpenStudy (anonymous):

Add all those up and subtract $15,000 to get: $1166.28 total gain.

OpenStudy (anonymous):

THANK YOU SOOOOOOO MUCH :)

OpenStudy (anonymous):

Any time!